During 20XX Artic Company completed its first year of operations: Below are selected data: Chapter 21 Homework Question 1: (1.0 point) Required: Using the variable costing approach, calculate the unit product cost. When entering the answers in Blackboard omit $ signs. Chapter 21 Homework Question 2: (1.0 point) Required: Using the variable costing approach, calculate the Income from operations. When entering the answers in Blackboard omit $ signs. Chapter 21 Homework Question 3: (1.0 point) Required: Using the absorption costing approach, calculate the unit product cost. When entering the answers in Blackboard omit $ signs. Chapter 21 Homework Question 4: (1.0 point) Required: Using the absorption costing approach, calculate the Income from operations. When entering the answers in Blackboard omit $ signs. Sales (12,000 units) $144,000 Production costs (15,000 units): Direct materials $12,000 Direct labor 24,000 Variable factory overhead 28,500 Fixed factory overhead 30,000 Total production costs 94,500 Operating expenses: Variable operating expenses $15,000 Fixed operating expenses 20,000 Total operating expenses 35,000
Answer #1

Part 1

Unit product cost (variable costing)

$4.30

Total variable production cost = Direct materials + Direct labor + Variable factory overhead = 12000+24000+28500 = 64500

Unit product cost = total variable production cost / total units produced = 64500/15000 = $4.30

Part 2

Income from operations (variable costing)

$30400

Sales

144000

Less: variable cost

Variable cost of goods sold (12000*4.30)

51600

Variable operating expenses (15000/15000*12000)

12000

63600

Contribution margin

$80400

Less: fixed costs

Fixed factory overhead

30000

Fixed operating expenses

20000

50000

Income from operations

$30400

Part 3

Unit product cost (fixed costing)

$6.30

Total Fixed production cost = Direct materials + Direct labor + Variable factory overhead + fixed factory overhead = 12000+24000+28500+ 30000 = 94500

Part 4

Income from operations (fixed costing)

$29200

Sales

144000

Less: cost of goods sold (6.30*12000)

75600

Gross profit

61200

Less: operating expenses

Variable operating expenses (15000/15000*12000)

12000

Fixed operating expenses

20000

32000

Income from operations

$29200